Well, no sooner is a book printed that you look at it and find all sorts of stupid errors that you should have caught before. Oh well, here are a few that I have noticed. If you find some more, let me know. Minor Errors

Page 37, line 3 from bottom: It is a universal R^{m} bundle.

Page 45: In the last displayed formula for φ_{i} the i-th coordinate should be y_{i} instead of x_{i}.

Page 122: It is stated that Z^{A_{S}} is a vector space, which of course it isn’t. I think in an earlier version, it was a vector space but was later changed to be the integer lattice instead. But somehow not all the text was changed accordingly.

Page 168: In the proof of Theorem 5.3.2 the first paragraph, one of the induction hypotheses is stated as q_{i}(V_{i}) is an algebraic set. This often could not be so, for example if W did not have constant local dimension. This hypothesis should be deleted, but in the case dim V_{i} < dim V_{j} for all i < j the extra induction hypothesis should be modified to ⋃_{j≤i} q_{j}(V_{j}) ⊃ Sing(⋃_{j≤i+1} q_{j}(V_{j})) for i = 0, …, n − 2. With these modifications, the proof becomes correct.